OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 2 ?

OpenStudy (anonymous):

The first step is to remember that log a - log b = log (a/b)

OpenStudy (anonymous):

And the second step is to exponentiate. What is the base of the logarithms. Is it assumed to be 10?

OpenStudy (anonymous):

ok so that would be log2 x/4

OpenStudy (anonymous):

the base is 2

OpenStudy (anonymous):

Oh.. yes. So then you have x/4 = 2^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but what next